OpenStudy (loser66):

{X_n} converges or diverges? \(X_n =\dfrac{(-1)^n}{n+1}\) Help me, please

OpenStudy (gorv):

what is n ??? n varies from???

OpenStudy (loser66):

\(n\in N\)

OpenStudy (gorv):

loll means 1 to infinty ??

OpenStudy (loser66):

yes

ganeshie8 (ganeshie8):

\[\large -1 \le (-1)^n \le 1\]

ganeshie8 (ganeshie8):

\[\dfrac{-1}{n+1} \le \dfrac{ (-1)^n}{n+1} \le \dfrac{ 1}{n+1}\]

ganeshie8 (ganeshie8):

take the limit through out the inequality

OpenStudy (gorv):

yeah but -1^n will remail 1 or -1 only not between -1 and 1

ganeshie8 (ganeshie8):

we're setting up for using squeeze theorem here @gorv

OpenStudy (loser66):

It shows lim = 0--> it converges to 0 I confused at upper and lower limits where the far left, 0 is upper limit and the far right , 0 is upper limit, also 

OpenStudy (gorv):

yeah that is right

ganeshie8 (ganeshie8):

since both upper and lower limits are 0, it converges to 0

OpenStudy (loser66):